OpenStudy (anonymous):

Find P(A and B) P(A) = 3/4, P(B given A) = 5/8

OpenStudy (anonymous):

Is this a triangle?

OpenStudy (anonymous):

P(A and B) = P(A) * P(B given A) = 3/4 * 5/8 = 15/32